OpenStudy (anonymous):

how do i simplify 9p - 2n + 4p + 2n

OpenStudy (anonymous):

Hello, Welcome to Openstudy!

OpenStudy (anonymous):

Combine like terms .

OpenStudy (anonymous):

Combine like terms

OpenStudy (anonymous):

n and p in this case

OpenStudy (anonymous):

9p-2n+4p+2n=...

OpenStudy (anonymous):

9p+4p_2n+2n=

OpenStudy (anonymous):

9p+4p+0=

OpenStudy (shamil98):

\[\huge 9p + 4p - 2n + 2n\] think of it as (9+4)p + (2-2)n

OpenStudy (anonymous):

9p+4p=

OpenStudy (shamil98):

what do you get?

OpenStudy (anonymous):

13p

OpenStudy (anonymous):

i got 13p?

OpenStudy (anonymous):

Plzzz. do not confuse this young girl with parentheses @shamil98 !!!!!

OpenStudy (anonymous):

very good @daisyfuentes71 !

OpenStudy (shamil98):

yes, 13p + 0n which is simply 13p.

OpenStudy (anonymous):

thank you for your help! @_CALCULASMASTER_

OpenStudy (anonymous):

yup np